# Inequality Restricted Minimum Density Power Divergence Estimation in Panel Count Data

### v1

@@ 7 Conclusion
National Health and Nutrition Examination Survey (NHANES); NHANES 2017-March 2020 Pre-pandemic Questionnaire Data -
https://www.cdc.gov/nchs/nhanes.
@@ Acknowledgements
The authors extend their gratitude to the anonymous reviewers, the associate editor, and the editor for their valuable and constructive suggestions, which have greatly improved the quality of this paper.
@@
Appendix
Proof of Theorem (a)
According to Hanson 1965, the demonstration of the theorem unfolds as described below.
@@
$g_{m}(\bm{\theta})=$ $\frac{\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})}{m(1+\gamma)}$ (A.1)
We adhere to the proof outline in Aitchison and Silvey 1958. For $\bm{\theta}_{0}\in\Theta_{0}$, the true parameter value, define $U_{\alpha}$ as the set $\left\{\bm{\theta}:\left\|\bm{\theta}-\bm{\theta}_{0}\right\|\leq\alpha\right\}$. There exists a constant $\delta>0$ such that $U_{\delta}\in\omega$, where $\omega$ is specified in regulatory condition, (R2). We aim to show that, under the constraint $\bm{h}\left(\bm{\theta}\right)\geqq\mathbf{0}_{r}$, the estimating equation for the density power divergence possesses a root in $U_{\delta}$ as $m\to\infty$.
@@
$m\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})+\bm{\lambda}\bm{H}\left(\bm{\theta}\right)=\mathbf{0}_{p}\$ (A.2)
$\bm{h}\left(\bm{\theta}\right)\geqq\mathbf{0}_{r}\$ (A.3)
$\bm{\lambda}^{\prime}\bm{h}\left(\bm{\theta}\right)=\mathbf{0}\$ (A.4)
$\bm{\lambda}\geqq\mathbf{0}\$ (A.5)
@@
$+\bm{\lambda}\bm{H}\left(\bm{\theta}_{0}\right)+\bm{\nu}_{m}^{(2)}(\bm{\theta})=\mathbf{0}_{p}$ (A.6)
@@
(i) The vector $\bm{\nu}_{m}^{(1)}(\bm{\theta})$ has its $c^{th}$ element given by $\frac{1}{2}\left(\bm{\theta}-\bm{\theta}_{0}\right)^{T}\bm{G}_{c}\left(\bm{\theta}-\bm{\theta}_{0}\right)$, where $\bm{G}_{c}$ is a matrix. The $(u,v)^{th}$ component of $\bm{G}_{c}$ is defined as
$\partial^{3}g_{m}(\bm{\theta})/\left.\partial\bm{\theta}_{c}\partial\bm{\theta}_{u}\partial\bm{\theta}_{v}\right|_{\bm{\theta}=\bm{\theta}_{0}^{(c,1)}}$, with indices $u,v,c=1,2,\cdots,p$. Additionally, the norm condition $\left\|\bm{\theta}^{(c,1)}-\bm{\theta}_{0}\right\|\leq\left\|\bm{\theta}-\bm{\theta}_{0}\right\|$ holds.
(ii) The vector $\bm{\nu}_{m}^{(2)}(\bm{\theta})$ has its $c^{th}$ element given by $\frac{1}{2}\left(\bm{\theta}-\bm{\theta}_{0}\right)^{T}\bm{W}_{c}\left(\bm{\theta}-\bm{\theta}_{0}\right)$, where $\bm{W}_{c}$ is a matrix. The $(u,v)^{th}$ component of $\bm{W}_{c}$ is defined as
$\partial^{2}\bm{h}_{m}(\bm{\theta})/\left.\partial\bm{\theta}_{i}\partial\bm{\theta}_{j}\right|_{\bm{\theta}=\bm{\theta}_{0}^{(c,2)}}$, with indices $u,v=1,2,\cdots,p$, and $c=1,2,\cdots,r$. Further, the norm condition $\left\|\bm{\theta}^{(c,2)}-\bm{\theta}_{0}\right\|\leq\left\|\bm{\theta}-\bm{\theta}_{0}\right\|$ holds.
@@
(i) $\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\rightarrow\mathbf{0}_{p}$.
(ii) $\left.\frac{\partial^{2}}{\partial\bm{\theta}^{T}\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\rightarrow\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)$.
(iii) The elements of $(1/m)\bm{G}_{c}$ for $c=1,2,\cdots,r$, remain bounded in $\bm{\theta}\in U_{\delta}$ as $m\rightarrow\infty$.
(iv) The elements of $(1/m)\bm{W}_{c}$ for $c=1,2,\cdots,r$, remain bounded in $\bm{\theta}\in U_{\delta}$ as $m\rightarrow\infty$.
@@
$\bm{J}_{\gamma}(\bm{\theta_{0}})\left(\bm{\theta}-\bm{\theta}_{0}\right)+\frac{\bm{\lambda}}{m}\bm{H}\left(\bm{\theta}_{0}\right)+\delta^{2}\bm{\nu}_{m}^{(3)}(\bm{\theta})=\mathbf{0}_{p}$ (A.7)
@@
Next, it will be shown that the system (A.2) - (A.5) has a solution ${\widehat{\bm{\theta}}}$, ${\widehat{\bm{\lambda}}}$, where ${\widehat{\bm{\theta}}}\in U_{\delta}$.
Since $\bm{\theta}_{0}$ satisfies the constraint, (A.3) we may write,
$\bm{h_{i}}\left(\bm{\theta_{0}}\right)=\mathbf{0},\quad i=1,2,\cdots,s$ (A.8)
@@
$\bm{h_{j}}\left(\bm{\theta_{0}}\right)-\bm{v_{j,0}}=\mathbf{0},\quad j=s+1,\cdots,r$ (A.9)
for some value of $s\leqq r$, where $\bm{v_{j,0}}>\mathbf{0},$ $j=s+1,\cdots,r$. Let $\bm{u_{0}}=\left(\bm{u_{s+1,0}}\right.,\left.\bm{u_{s+2,0}},\cdots,\bm{u_{r,0}}\right)$. If $\bm{\lambda_{0}}=\left(\bm{\lambda_{1,0}},\bm{\lambda_{2,0}},\cdots,\bm{\lambda_{r,0}}\right)$ is accordingly defined in such a way that
$\bm{\lambda_{k,0}}\geqq\mathbf{0},\quad k=1,2,\cdots,s,$ (A.10)
@@
$\bm{\lambda_{l,0}}=\mathbf{0},\quad l=s+1,\cdots,r$ (A.11)
@@
$m\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})+\bm{\lambda^{s}}\bm{H_{s,\theta}}=\mathbf{0}$ (A.12)
$\bm{\lambda_{k}}\geqq\mathbf{0},\quad k=1,2,\cdots,s$ (A.13)
$\bm{\lambda_{l}}=\mathbf{0},\quad k=s+1,s+2,\cdots,r$ (A.14)
$\bm{u_{j}}>\mathbf{0},\quad j=s+1,\cdots,r$ (A.15)
where $\bm{H_{s;\bm{\theta}_{0}}}$ is the $p\times s$ matrix $\left[\frac{\partial\bm{h}_{i}(\bm{\theta})}{\partial\bm{\theta}_{k}}\right],i=1,2,\cdots,s$;  $k=1,2,\cdots,p$ and $\bm{\lambda}^{(s)}$ is the vector of dimensions $s$ whose elements are the first $s$ elements of $\bm{\lambda}$. Let $\bm{u}=\left(\bm{u}_{s+1},\bm{u}_{s+2},\cdots,\bm{u}_{r}\right)$. Let $\bm{h}^{(s)}(\bm{\theta})=\left(\bm{h}_{1}(\bm{\theta}),\bm{h}_{2}(\bm{\theta}),\cdots,\bm{h}_{s}(\bm{\theta})\right)$. Expanding $\bm{h}^{(s)}(\bm{\theta})$ about $\bm{\theta}_{0}$, we have
$\bm{h^{(s)}}(\bm{\theta})=\bm{h^{(s)}}\left(\bm{\theta_{0}}\right)+\bm{H_{s;\bm{\theta_{0}}}}\left(\bm{\theta}-\bm{\theta_{0}}\right)+\bm{\delta^{2}}\bm{\nu}_{m}^{(4)}(\bm{\theta})$ (A.16)
@@
$\bm{h^{\left(s^{*}\right)}}(\bm{\theta})=\bm{h^{\left(s^{*}\right)}}\left(\bm{\theta_{0}}\right)+\bm{H_{s^{*};\bm{\theta_{0}}}}^{T}\left(\bm{\theta}-\bm{\theta_{0}}\right)+\bm{\delta^{2}}\bm{\nu}_{m}^{(5)}(\bm{\theta}),\$ (A.17)
where $\bm{H_{s^{*};\bm{\theta_{0}}}}$ is the $p\times(r-s)$ matrix $\left[\frac{\partial\bm{h}_{j}(\bm{\theta})}{\partial\bm{\theta}_{k}}\right],j=s+1,\cdots,r;\quad k=1,2,\cdots,p$ and $\left\|\bm{\nu}_{m}^{(5)}(\bm{\theta})\right\|$ remains bounded for $\bm{\theta}\in U_{\delta}$.
@@
$\bm{J}_{\gamma}(\bm{\theta}_{0})\left(\bm{\theta}-\bm{\theta}_{0}\right)+\frac{\bm{\lambda^{(s)}}}{m}\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}\bm{\nu}_{m}^{(4)}(\bm{\theta})=\mathbf{0}_{p}\$ (A.18)
@@
$\bm{H_{s;\bm{\theta_{0}}}}^{T}\left(\bm{\theta}-\bm{\theta}_{0}\right)+\left(\frac{\bm{\lambda^{(s)}}}{m}\right)\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})=\mathbf{0}_{p}\$ (A.19)
@@
$\left(\frac{\bm{\lambda^{(s)}}}{m}\right)\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})+\bm{\nu}_{m}^{(5)}(\bm{\theta})]=\mathbf{0}_{p}$ (A.20)
As $(\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}})$ is non-singular, its inverse remains bounded for $\bm{\theta}\in U_{\delta}$. Consequently, equation (A.20) can be solve for $\bm{\lambda^{(s)}}$ in terms of $\bm{\theta}.$ Let it be $\bm{\hat{\lambda}^{(s)}}$, then
$\frac{\bm{\hat{\lambda}^{(s)}}}{m}=[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}]^{-1}\delta^{2}[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})+\bm{\nu}_{m}^{(5)}(\bm{\theta})]$ (A.21)
@@
$\bm{J}_{\gamma}\left(\bm{\theta}-\bm{\theta}_{0}\right)+\delta^{2}\bm{\nu}^{*}(\bm{\theta})=\mathbf{0}$ (A.22)
@@
Proof of Theorem (b)
This proof closely follows the approach taken by Sen et al. 2010. Differentiating both sides of equation (A.1) with respect to $\bm{\theta}$, we obtain
$\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})=\left.\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})f^{\gamma+1}(\bm{n})\right.-\left.\frac{1}{m}\sum_{i=1}^{m}\bm{u}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}})\right.$ (A.23)

### v4

@@ 7 Conclusion
National Health and Nutrition Examination Survey (NHANES); NHANES 2017-March 2020 Pre-pandemic Questionnaire Data - https://www.cdc.gov/nchs/nhanes.
@@ Acknowledgements
The authors extend their gratitude to the anonymous reviewers, the administrative editor, the associate editor, and the editor for their valuable and constructive suggestions, which have greatly improved the quality of this paper.
@@
Appendix A Proof of Theorem (a)
According to Hanson [42], the demonstration of the theorem unfolds as described below.
@@ Appendix A Proof of Theorem (a)
$g_{m}(\bm{\theta})=$ $\frac{\sum_{i=1}^{m}V_{\bm{\theta}}(\bm{N_{i}})}{m(1+\gamma)}.$ (A.1)
We adhere to the proof outline in Aitchison and Silvey [43]. For $\bm{\theta}_{0}\in\Theta_{0}$, the true parameter value, define $U_{\alpha}$ as the set $\left\{\bm{\theta}:\left\|\bm{\theta}-\bm{\theta}_{0}\right\|\leq\alpha\right\}$. There exists a constant $\delta>0$ such that $U_{\delta}\in\omega$, where $\omega$ is specified in regulatory condition, (R2). We aim to show that, under the constraint $\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r}$, the estimating equation for the density power divergence possesses a root in $U_{\delta}$ as $m\to\infty$.
@@
    For every $\bm{\theta}\in U_{a}$, the partial derivatives $\frac{\partial\bm{h}_{k}(\bm{\theta})}{\partial\bm{\theta}_{i}}$, for $i=1,2,\dots,p$ and $k=1,2,\dots,r$, exist and are continuous functions of $\bm{\theta}$.
@@
    For every $\bm{\theta}\in U_{\alpha}$, the second-order partial derivatives $\frac{\partial^{2}\bm{h}_{k}(\bm{\theta})}{\partial\bm{\theta}_{i}\partial\bm{\theta}_{j}}$, for $i,j=1,2,\dots,p$ and $k=1,2,\dots,r$, exist and satisfy $\left|\frac{\partial^{2}\bm{h}_{k}(\bm{\theta})}{\partial\bm{\theta}_{i}\partial\bm{\theta}_{j}}\right|<2\kappa_{2}$, where $\kappa_{2}$ is a given constant.
@@
    The $p\times r$ matrix $\bm{H}_{\bm{\theta}}=\left[\frac{\partial\bm{h}_{j}(\bm{\theta})}{\partial\bm{\theta}_{i}}\right]$, evaluated at $\bm{\theta}=\bm{\theta}_{0}$, has maximum rank. If $r>p$, then at most $p$ components of $\bm{h}$ can vanish at each point of $U_{\alpha}$. For the components $\bm{h}_{k}$, say for $k=1,2,\dots,s$, that vanish at some point in $U_{\alpha}$, the $p\times s$ matrix $\left[\frac{\partial\bm{h}_{k}(\bm{\theta})}{\partial\bm{\theta}_{i}}\right]$ also has maximum rank.
@@
$m\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})+\bm{\lambda}\bm{H}\left(\bm{\theta}\right)=\mathbf{0}_{p},\\$ (A.2)
$\bm{h}\left(\bm{\theta}\right)\geq\mathbf{0}_{r},\\$ (A.3)
$\bm{\lambda}^{\prime}\bm{h}\left(\bm{\theta}\right)=\mathbf{0},\\$ (A.4)
$\bm{\lambda}\geq\mathbf{0},\\$ (A.5)
@@
$+\bm{\lambda}\bm{H}\left(\bm{\theta}_{0}\right)+\bm{\nu}_{m}^{(2)}(\bm{\theta})=\mathbf{0}_{p},$ (A.6)
@@
  • $\bm{(i)}$
    The vector $\bm{\nu}_{m}^{(1)}(\bm{\theta})$ has its $c^{th}$ element given by $\frac{1}{2}\left(\bm{\theta}-\bm{\theta}_{0}\right)^{T}\bm{G}_{c}\left(\bm{\theta}-\bm{\theta}_{0}\right)$, where $\bm{G}_{c}$ is a matrix. The $(u,v)^{th}$ component of $\bm{G}_{c}$ is defined as
    $\partial^{3}g_{m}(\bm{\theta})/\left.\partial\bm{\theta}_{c}\partial\bm{\theta}_{u}\partial\bm{\theta}_{v}\right|_{\bm{\theta}=\bm{\theta}_{0}^{(c,1)}}$, with indices $u,v,c=1,2,\dots,p$. Additionally, the norm condition $\left\|\bm{\theta}^{(c,1)}-\bm{\theta}_{0}\right\|\leq\left\|\bm{\theta}-\bm{\theta}_{0}\right\|$ holds.
  • $\bm{(ii)}$
    The vector $\bm{\nu}_{m}^{(2)}(\bm{\theta})$ has its $c^{th}$ element given by $\frac{1}{2}\left(\bm{\theta}-\bm{\theta}_{0}\right)^{T}\bm{W}_{c}\left(\bm{\theta}-\bm{\theta}_{0}\right)$, where $\bm{W}_{c}$ is a matrix. The $(u,v)^{th}$ component of $\bm{W}_{c}$ is defined as
    $\partial^{2}\bm{h}_{m}(\bm{\theta})/\left.\partial\bm{\theta}_{i}\partial\bm{\theta}_{j}\right|_{\bm{\theta}=\bm{\theta}_{0}^{(c,2)}}$, with indices $u,v=1,2,\dots,p$, and $c=1,2,\dots,r$. Further, the norm condition $\left\|\bm{\theta}^{(c,2)}-\bm{\theta}_{0}\right\|\leq\left\|\bm{\theta}-\bm{\theta}_{0}\right\|$ holds.
@@
  • $\bm{(i)}$
    $\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\rightarrow\mathbf{0}_{p}$.
  • $\bm{(ii)}$
    $\left.\frac{\partial^{2}}{\partial\bm{\theta}^{T}\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\rightarrow\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)$.
  • $\bm{(ii)}$
    The elements of $(1/m)\bm{G}_{c}$ for $c=1,2,\dots,r$, remain bounded in $\bm{\theta}\in U_{\delta}$ as $m\rightarrow\infty$.
  • $\bm{(ii)}$
    The elements of $(1/m)\bm{W}_{c}$ for $c=1,2,\dots,r$, remain bounded in $\bm{\theta}\in U_{\delta}$ as $m\rightarrow\infty$.
@@
$\bm{J}_{\gamma}(\bm{\theta_{0}})\left(\bm{\theta}-\bm{\theta}_{0}\right)+\frac{\bm{\lambda}}{m}\bm{H}\left(\bm{\theta}_{0}\right)+\delta^{2}\bm{\nu}_{m}^{(3)}(\bm{\theta})=\mathbf{0}_{p},$ (A.7)
@@
Next, it will be shown that the system (A.2) - (A.5) has a solution $^{\widehat{\bm{\theta}}}$, $^{\widehat{\bm{\lambda}}}$, where ${\widehat{\bm{\theta}}}\in U_{\delta}$.
Since $\bm{\theta}_{0}$ satisfies the constraint, (A.3) we may write
$\bm{h_{i}}\left(\bm{\theta_{0}}\right)=\mathbf{0},\quad i=1,2,\dots,s,$ (A.8)
@@
$\bm{h_{j}}\left(\bm{\theta_{0}}\right)-\bm{u_{j,0}}=\mathbf{0},\quad j=s+1,\dots,r.$ (A.9)
for some value of $s\leq r$, where $\bm{u_{j,0}}>\mathbf{0},$ $j=s+1,\dots,r$. Let $\bm{u_{0}}=\left(\bm{u_{s+1,0}}\right.,\left.\bm{u_{s+2,0}},\dots,\bm{u_{r,0}}\right)$. If $\bm{\lambda_{0}}=\left(\bm{\lambda_{1,0}},\bm{\lambda_{2,0}},\dots,\bm{\lambda_{r,0}}\right)$ is accordingly defined in such a way that
$\bm{\lambda_{k,0}}\geq\mathbf{0},\quad k=1,2,\dots,s,$ (A.10)
@@
$\bm{\lambda_{l,0}}=\mathbf{0},\quad l=s+1,\dots,r,$ (A.11)
@@
$m\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})+\bm{\lambda^{\textbf{(s)}}}\bm{H_{s,\theta}}=\mathbf{0},$ (A.12)
$\bm{\lambda_{k}}\geq\mathbf{0},\quad k=1,2,\dots,s,$ (A.13)
$\bm{\lambda_{l}}=\mathbf{0},\quad k=s+1,s+2,\dots,r,$ (A.14)
$\bm{u_{j}}>\mathbf{0},\quad j=s+1,\dots,r,$ (A.15)
where $\bm{H_{s;\bm{\theta}}}$ is the $p\times s$ matrix $\left[\frac{\partial\bm{h}_{i}(\bm{\theta})}{\partial\bm{\theta}_{k}}\right],i=1,2,\dots,s$;  $k=1,2,\dots,p$ and $\bm{\lambda}^{\textbf{(s)}}$ is the vector of dimensions $s$ whose elements are the first $s$ elements of $\bm{\lambda}$. Let $\bm{u}=\left(\bm{u}_{s+1},\bm{u}_{s+2},\dots,\bm{u}_{r}\right)$. Let $\bm{h}^{\textbf{(s)}}(\bm{\theta})=\left(\bm{h}_{1}(\bm{\theta}),\bm{h}_{2}(\bm{\theta}),\dots,\bm{h}_{s}(\bm{\theta})\right)$. Expanding $\bm{h}^{\textbf{(s)}}(\bm{\theta})$ about $\bm{\theta}_{0}$, we have
$\bm{h^{\textbf{(s)}}}(\bm{\theta})=\bm{h^{\textbf{(s)}}}\left(\bm{\theta_{0}}\right)+\bm{H_{s;\bm{\theta_{0}}}}^{T}\left(\bm{\theta}-\bm{\theta_{0}}\right)+\bm{\delta^{2}}\bm{\nu}_{m}^{(4)}(\bm{\theta}),$ (A.16)
@@
$\bm{h^{\left(\textbf{s}^{*}\right)}}(\bm{\theta})=\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)+\bm{H_{s^{*};\bm{\theta_{0}}}}^{T}\left(\bm{\theta}-\bm{\theta_{0}}\right)+\bm{\delta^{2}}\bm{\nu}_{m}^{(5)}(\bm{\theta}),\\$ (A.17)
where $\bm{H_{s^{*};\bm{\theta_{0}}}}$ is the $p\times(r-s)$ matrix $\left[\frac{\partial\bm{h}_{j}(\bm{\theta})}{\partial\bm{\theta}_{k}}\right],j=s+1,\dots,r;\quad k=1,2,\dots,p$ and $\left\|\bm{\nu}_{m}^{(5)}(\bm{\theta})\right\|$ remains bounded for $\bm{\theta}\in U_{\delta}$.
@@
$\bm{J}_{\gamma}\left(\bm{\theta}-\bm{\theta}_{0}\right)+\frac{\bm{\lambda^{(s)}}}{m}\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}\bm{\nu}_{m}^{(4)}(\bm{\theta})=\mathbf{0}_{p}.\\$ (A.18)
@@
$\bm{H_{s;\bm{\theta_{0}}}}^{T}\left(\bm{\theta}-\bm{\theta}_{0}\right)+\left(\frac{\bm{\lambda^{\textbf{(s)}}}}{m}\right)\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})=\mathbf{0}_{p}.\\$ (A.19)
@@
$\left(\frac{\bm{\lambda^{\textbf{(s)}}}}{m}\right)\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}+\delta^{2}[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})+\bm{\nu}_{m}^{(5)}(\bm{\theta})]=\mathbf{0}_{p}.$ (A.20)
As $(\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}})$ is non-singular, its inverse remains bounded for $\bm{\theta}\in U_{\delta}$. Consequently, equation (A.20) can be solved for $\bm{\lambda^{\textbf{(s)}}}$ in terms of $\bm{\theta}.$ Let it be $\bm{\hat{\lambda}^{\textbf{(s)}}}$, then
$\frac{\bm{\hat{\lambda}^{\textbf{(s)}}}}{m}=[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{H_{s;\bm{\theta_{0}}}}]^{-1}\delta^{2}[\bm{H_{s;\bm{\theta_{0}}}}^{T}\bm{J^{-1}_{\gamma}}(\bm{\theta}_{0})\bm{\nu}_{m}^{(4)}(\bm{\theta})+\bm{\nu}_{m}^{(5)}(\bm{\theta})].$ (A.21)
@@
$\bm{J}_{\gamma}\left(\bm{\theta}-\bm{\theta}_{0}\right)+\delta^{2}\bm{\nu}^{*}(\bm{\theta})=\mathbf{0},$ (A.22)
@@
Appendix B Proof of Theorem (b)
This proof closely follows the approach taken by Sen et al. [44]. Differentiating both sides of equation (A.1) with respect to $\bm{\theta}$, we obtain
$\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})=\left.\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})f^{\gamma+1}(\bm{n})\right.-\left.\frac{1}{m}\sum_{i=1}^{m}\bm{u}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}})\right$., (B.1)